# TokaMark: A Comprehensive Benchmark for MAST Tokamak Plasma Models

### v1

@@ Abstract.
Development and operation of commercially viable fusion energy reactors such as tokamaks require accurate predictions of plasma dynamics from sparse, noisy, and incomplete sensors readings. The complexity of the underlying physics and the heterogeneity of experimental data pose formidable challenges for conventional numerical methods, while simultaneously highlights the promise of modern data-native AI approaches. A major obstacle in realizing this potential is, however, the lack of curated, openly available datasets and standardized benchmarks. Existing fusion datasets are scarce, fragmented across institutions, facility-specific, and inconsistently annotated, which limits reproducibility and prevents a fair and scalable comparison of AI approaches. In this paper, we introduce TokaMark, a structured benchmark to evaluate AI models on real experimental data collected from the Mega Ampere Spherical Tokamak (MAST). TokaMark provides a comprehensive suite of tools designed to (i) unify access to multi-modal heterogeneous fusion data (ii) harmonize formats, metadata, temporal alignment and evaluation protocols to enable consistent cross-model and cross-task comparisons. The benchmark includes a curated list of 14 tasks spanning a range of physical mechanisms, exploiting a variety of diagnostics and covering multiple target use cases. A baseline model is provided to facilitate transparent comparison and validation within a unified framework. By establishing a unified benchmark for both the fusion and AI-for-science communities, TokaMark  aims to accelerate progress in data-driven plasma AI modeling, contributing to the broader goal of achieving sustainable and stable fusion energy. The benchmark, documentation, and tooling will be fully open sourced upon acceptance to encourage community adoption and contribution.
AI for science, Fusion benchmark, Data-driven plasma modeling, Multi-modal learning, MAST tokamak
copyright: nonecopyright: nonecopyright: noneccs: Hardware Renewable energyccs: Software and its engineering Software libraries and repositoriesccs: Software and its engineering Open source modelccs: Information systems Extraction, transformation and loadingccs: Information systems Evaluation of retrieval results
@@ 1. Introduction
Nuclear fusion power is being explored as a potential long-term energy source with a unique combination of benefits. It offers the prospect of a carbon-neutral energy supply with abundant fuel and significant safety advantages over nuclear fission. However commercially viable fusion demands stable, sustained reactions that produce more power than the reactor consumes – a goal made difficult by the extreme physical conditions of thermonuclear confinement under which fusion must operate (Donné et al., 2025).
Fusion reactors such as tokamaks aim to reproduce on Earth the same process that powers the Sun, confining a plasma hotter than 100 million degrees Celsius using strong rapidly evolving magnetic fields because no material surface can survive direct contact with such plasma. This environment forces all measurements to be non-invasive and indirect, only partially inferring plasma state while the underlying dynamics evolve on microsecond-to-millisecond timescales. In this work, we dedicate our attention to the problem of modeling fusion plasma dynamics in tokamaks – one of the central challenges in fusion research. This problem encompasses a wide set of predictive tasks, including plasma shape and equilibrium inference, transport and profile evolution, and forecasting of MHD activity and disruptions.
@@ 1.1. AI for fusion plasma modeling
Traditional tokamak plasma modeling approaches rely either on first‑principles, physics‑driven solvers or on complex system identification pipelines. Although the governing physics is well understood, it is expressed through complex nonlinear partial differential equations whose key parameters are often unmeasured, only partially observable, or must be inferred indirectly from diagnostics. As a result, classical methods tend to be computationally intensive, brittle across operating regimes, and difficult to apply in real time. These challenges are further exacerbated by the nature of the experimental data itself, which reflects the extreme environment and sensing limitations of tokamak operation.
@@
Tokamaks deploy a broad suite of heterogeneous diagnostics – magnetics, optical and X‑ray emission, microwave interferometry, and more – mounted on or behind the reactor walls and engineered to withstand heat, radiation, and electromagnetic stress. These sensors operate at widely different sampling rates, spatial resolutions, and noise characteristics producing data that is inherently heterogeneous, incomplete, multi‑rate and noisy.
@@
Together, the complexity of the underlying physics and the heterogeneity of the data create an opportunity where modern AI methods can provide significant advantages. Data‑driven models are well suited to fusing multi‑modal diagnostics, learning latent plasma representations, and capturing nonlinear, multi‑scale temporal behavior. Unlike traditional solvers, AI models can operate directly on raw measurements, handle missing or asynchronous data, and produce accurate and efficient surrogates. This positions AI as a promising complement to physics‑based modeling for understanding, predicting, and ultimately controlling fusion plasma.
@@
Previous work applying AI to fusion plasma problems have demonstrated promising results across variety of narrowly defined tasks, ranging from plasma shape reconstruction and profile forecasting to actuator optimization and disruptions prediction. However, most of those efforts use bespoke pipelines tailored to a small set of diagnostics, single device and single scientific objective, relying heavily on task-specific feature engineering and handcrafted labeling procedures. While successful within their respective scopes, these approaches typically optimize models for individual tasks and limit reuse across the experimental lifecycle. Much more work is needed to move beyond isolated point solutions toward broad, interoperable models capable of understanding fusion plasmas in a comprehensive way.
@@
Progress toward either specialized or generalist plasma models is hampered by the absence of open, standardized benchmarks. Fusion datasets remain fragmented across institutions, locked behind proprietary interfaces, or stored in domain‑specific formats that are difficult for Machine Learning researchers to access or interpret (Strand et al., 2022). Without unified task definitions, metrics, or evaluation protocols, it becomes impossible to provide fair comparison of methods and to measure progress systematically. A benchmark is therefore needed to frame the plasma modeling problems in a broader and more structured way, enabling cross‑comparison across algorithms, reproducibility across labs, and accessibility for researchers both inside and outside the fusion community.
At the same time, fusion diagnostic data provides unique set of challenge going beyond those found in standard datasets from more established areas of AI. The data is inherently multi‑modal, multi‑rate, and multi‑dimensional, with wide variations in signal quality, noise characteristics, and information content (see Figure 1 for an illustrative example). Many diagnostics operate asynchronously, contain gaps, or suffer partial failures; others produce high‑dimensional structured outputs such as flux maps or spectrograms. This complexity demands dedicated modeling approaches that can handle missing data, fuse heterogeneous signals, different modalities, and reason across fast and slow timescales – capabilities that are difficult to stress‑test without a standardized benchmark.
@@ 1.2. TokaMark Benchmark Overview
To fill the need of comprehensive fusion plasma benchmarks we introduce TokaMark  the first large, open benchmark for evaluating AI models trained on real fusion diagnostics, built from recently released FAIR-MAST data (Jackson et al., 2025, 2024).
Data. To the best of our knowledge, FAIR-MAST is the only openly available dataset of real tokamak diagnostics. Recent releases (Jackson et al., 2025, 2024) curate a collection of real experiments and corresponding diagnostics measurements from MAST tokamak. From FAIR-MAST, we select 39 signals across heterogeneous modalities, harmonize metadata and build standardized loaders.
Tasks. For TokaMark we defined a diverse suite of 14 downstream tasks designed to probe core capabilities required of AI models for fusion plasmas: (i) representation learning from heterogeneous diagnostics; (ii) temporal reasoning across fast and slow timescales; (iii) robustness to incomplete state information; and (iv) generalization across operating regimes. Rather than optimizing for a single downstream objective, the tasks, organized into four groups, span a cascade of physical processes – from fast magnetic response to slower, transport-driven evolution and long-horizon precursors of MHD activity – while remaining closely aligned with routine experimental workflows. Wherever possible, tasks minimize reliance on expert-labeled targets, supporting self-supervised and weakly supervised formulations and enabling systematic evaluation of transferable plasma representations.
Evaluation. We introduce a hierarchical evaluation protocol aligned with the structure of FAIR-MAST and with scientific objectives. This hierarchy assesses both low-level prediction quality and high-level scientific meaning. For specialized models (single-task training), the hierarchy yields granular, signal-level diagnostics; for generalist models (FM-style pretraining across tasks), it provides modular comparisons across tasks and broad scientific objectives.
Baseline. Finally, we provide a strong yet accessible baseline: a multi-branch convolutional encoder–decoder architecture trained per each task that ingests heterogeneous inputs and predicts task-specific targets.
@@
  • Benchmark design. We define 14 tasks organized into four groups, with a hierarchical evaluation protocol, standardized windowing and error metrics.
  • Data packaging. We freeze a stable subset of FAIR-MAST, resolve schema inconsistencies, standardize metadata and units, and release it for reproducibility and long time compatibility.
  • Tools and API. We provide a Python package for task-specific data loading, processing and batching, masking and alignment utilities, and evaluation logic, integrated with PyTorch stack.
  • Baseline models. We release a multi-branch convolutional encoder–decoder baseline with reference configurations and training scripts, establishing reproducible baselines across all tasks (Seo et al., 2023) (Seo et al., 2024).
2. Preliminaries
To ground TokaMark benchmark in the reality of complex heterogeneous fusion data, we first introduce a signal taxonomy that organizes these measurements by function and modality, and allows us to provide a consistent vocabulary for describing inputs, targets, and tasks in the benchmark.
2.1. Data Taxonomy
@@
Magnetics Flux loops, pickup coils; saddle coils Sensor 5 kHz; 50 kHz Profile
Kinetics Thomson scattering; interferometer Sensor 0.2 kHz; 4 kHz Profile; time series
@@
MHD Precursors Mirnov coils Sensor 500 kHz Profile
Currents Poloidal Field Coil currents; Solenoid current, Plasma current Sensor 4 kHz Profile; Time series
@@
Equilibrium Shape parameters, $\beta$ metrics; plasma boundary; flux map Derived 0.2 kHz Time series; profile; video
@@
MAST is a spherically-shaped tokamak located in Culham, Oxfordshire (UK) and was operated by UKAEA and EURATOM from 1999 to 2013 (Sykes et al., 2001; Counsell et al., 2005; Meyer et al., 2009). MAST and tokamaks in general operate in short experimental cycles known as discharges or shots. The length of those cycles depends on the device size and in the case of MAST they are typically lasting around 2–3 seconds. Over its operational lifetime, MAST produced more than 30,000 shots, with each shot containing diagnostic signals measuring various properties of the plasma, including the magnetic field, plasma temperature, shape parameters, applied heating etc.
Recent works (Jackson et al., 2025, 2024) have created an open dataset of diagnostic data from a subset of the history of MAST. The FAIR-MAST dataset contains 11,573 shots from the last five experimental campaigns on MAST. In this work we utilized a total of 39 signals from FAIR-MAST in the design of our benchmark tasks.
@@
We organize all signals along several complementary axes, summarized in Table 1. First, we group signals by category, which encodes their physical semantics–that is, the physical quantities they represent and how they are used by MAST workflows. This information corresponds to the first two columns of Table 1. Note that the second column contains both individual signals and subcategories. These intermediate groupings are introduced for brevity: they consolidate multiple signals that are typically used together (at least within the proposed benchmark) and share the same structural and functional properties. For example, the shape parameters subcategory includes several attributes jointly describing plasma geometry. A full list of the signals used in this work is provided in the appendix in Table 4.
@@
Second, we distinguish signals by their origin. We define three classes: (i) sensors, corresponding to direct hardware measurements; (ii) actuators, representing controllable machine parameters used to steer plasma behavior; and (iii) derived signals, comprising quantities produced by reconstruction pipelines such as EFIT (e.g., shape parameters or flux maps).
@@
The third axis is frequency, reflecting the wide range of sampling rates across diagnostics, and spanning from 0.2 kHz up to 500 kHz. Finally, the fourth axis is modality, describing the structural form of each signal: (i) time series, represented as 1D tensors (scalar value over time); (ii) profiles, represented as 2D tensors (vector over time); and (iii) videos, represented as 3D tensors (image over time). These latter two axes are particularly important for guiding the design of AI model architectures and loss functions.
2.2. Tasks Structural Taxonomy

### v2

@@ Abstract.
Development and operation of commercially viable fusion energy reactors such as tokamaks require accurate predictions of plasma dynamics from sparse, noisy, and incomplete sensors readings. The complexity of the underlying physics and the heterogeneity of experimental data pose formidable challenges for conventional numerical methods, while simultaneously highlight the promise of modern data-native AI approaches. A major obstacle in realizing this potential is, however, the lack of curated, openly available datasets and standardized benchmarks. Existing fusion datasets are scarce, fragmented across institutions, facility-specific, and inconsistently annotated, which limits reproducibility and prevents a fair and scalable comparison of AI approaches. In this paper, we introduce TokaMark, a structured benchmark to evaluate AI models on real experimental data collected from the Mega Ampere Spherical Tokamak (MAST). TokaMark provides a comprehensive suite of tools designed to (i) unify access to multi-modal heterogeneous fusion data, and (ii) harmonize formats, metadata, temporal alignment and evaluation protocols to enable consistent cross-model and cross-task comparisons. The benchmark includes a curated list of 14 tasks spanning a range of physical mechanisms, exploiting a variety of diagnostics and covering multiple operational use cases. A baseline model is provided to facilitate transparent comparison and validation within a unified framework. By establishing a unified benchmark for both the fusion and AI-for-science communities, TokaMark  aims to accelerate progress in data-driven AI-based plasma modeling, contributing to the broader goal of achieving sustainable and stable fusion energy. The benchmark, documentation, and tooling will be fully open sourced upon acceptance to encourage community adoption and contribution.
AI for science, Data-driven modeling, Fusion benchmark, Plasma dynamics, Multi-modal learning, MAST tokamak
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: 32nd SIGKDD Conference on Knowledge Discovery and Data Mining, 2026; August 09–13, 2026; Jeju, Koreaisbn: XXX-X-XXXX-XXXX-X/2026/08copyright: nonecopyright: nonecopyright: noneccs: Applied computing Physical sciences and engineeringccs: Computing methodologies Knowledge representation and reasoningccs: Software and its engineering Software libraries and repositoriesccs: Software and its engineering Open source modelccs: Information systems Extraction, transformation and loading
@@ 1. Introduction
Nuclear fusion power is being explored as a potential long-term energy source with a unique combination of benefits. It offers the prospect of a carbon-neutral energy supply with abundant fuel and significant safety advantages over nuclear fission. However, commercially viable fusion demands stable, sustained operation that produces more energy than the power plant consumes—a goal made difficult by the extreme physical conditions of thermonuclear confinement under which fusion must operate (Donné et al., 2025).
Magnetic confinement fusion reactors such as tokamaks aim to reproduce the process that powers the Sun in a controlled environment, confining a plasma hotter than 100 million degrees Celsius through the use of strong magnetic fields because no material surface can survive direct contact with such plasma (Wesson, 1987; Freidberg, 2007). This environment forces all measurements to be non-invasive and indirect, only partially inferring plasma state while the underlying dynamics evolve on microsecond-to-millisecond timescales. In this work, we dedicate our attention to the problem of modeling fusion plasma dynamics in tokamaks—one of the central challenges in fusion research. This problem encompasses a wide set of predictive tasks, including plasma shape and equilibrium inference, transport and profile evolution, and forecasting of magnetohydrodynamics (MHD) activity and disruptions.
@@ 1.1. AI for fusion plasma modeling
Traditional approaches to tokamak plasma modeling are rooted in well-established first-principles descriptions of magnetized plasma dynamics. These descriptions are expressed through coupled, nonlinear systems of partial differential equations, whose numerical solutions often requires high-fidelity, multi-scale simulations. While such models are indispensable for predictive studies, their computational cost severely limits their routine use, both in exploring the full phenomenology of plasma behavior, and in systematically interrogating large experimental datasets. In particular, many physically relevant regimes remain difficult to characterize in detail because comprehensive parameter scans and high-fidelity simulations are prohibitively expensive.
The same computational burden complicates data-driven inference. Key parameters governing plasma behavior—such as transport coefficients, source terms, or stability-relevant profile features—are frequently unmeasured or only indirectly observable, and inferring them typically requires repeated forward simulations embedded within optimization or system identification loops. As a result, fitting models to experimental data becomes costly and brittle across operating regimes. These limitations make the direct use of first-principles solvers infeasible for real-time applications, where stringent latency constraints rule out iterative or high-fidelity numerical solutions. The challenge is further compounded by the experimental characteristics of tokamak data.
@@
Tokamaks deploy a broad suite of heterogeneous diagnostics–magnetics, optical and X‑ray emission, microwave interferometry, and more (Morris and the MAST team., 2002)—mounted on or behind the reactor walls and engineered to withstand extreme heat, radiation, and electromagnetic stress. These sensors operate at widely different sampling rates, spatial resolutions, and noise characteristics, producing data that is inherently heterogeneous, incomplete, multi‑rate, and noisy.
@@
Together, the complexity of the underlying physics and the heterogeneity of the data create an opportunity where modern AI methods can provide significant advantages and complement physics‑based modeling. Data‑driven models are well suited to fusing multi‑modal diagnostics, learning latent plasma representations, and capturing nonlinear, multi‑scale temporal behavior without explicit physical parameterization. Unlike traditional solvers, AI models can operate directly on raw measurements, handle missing or asynchronous data, and produce accurate and efficient surrogates. However, these advantages come with caveats: learned models may fail silently when operating outside their training distribution, lack guaranteed physical consistency, and are generally more difficult to interpret.
@@
Previous work applying AI to fusion plasma problems have demonstrated promising results across a variety of narrowly defined tasks, ranging from plasma shape reconstruction (Wan et al., 2023; Wai et al., 2022; Rossi et al., 2023) and profile forecasting (Wan et al., 2021; Abbate et al., 2021, 2023, 2025; Char et al., 2024; Kit et al., 2024; Wakatsuki et al., 2023) to actuator optimization (Wang et al., 2025, 2024; Yang et al., 2020; Schramm et al., 2024; Seo et al., 2024, 2021, 2022; Vega et al., 2022; Degrave et al., 2022; Abbate et al., 2023; Orozco et al., 2022) and disruptions prediction (Zhang et al., 2020; Zhu et al., 2023; Priyanka et al., 2024; Rea et al., 2018; Lucas et al., 2024; Montes, 2021; Montes et al., 2021; Churchill et al., 2020; Ferreira et al., 2020; Guo et al., 2021; Kates-Harbeck et al., 2019; Zhu et al., 2021b; De Vries et al., 2011; Zhu et al., 2021a; Aymerich et al., 2022). However, most of those efforts use bespoke pipelines tailored to a small set of diagnostics, single device and single scientific objective, relying heavily on task-specific feature engineering and handcrafted labeling procedures. While successful within their respective scopes, these approaches typically optimize models for individual tasks, which limits reuse across the experimental lifecycle. Adding to ongoing efforts (Dong et al., 2025; Yang et al., 2025), more work is needed to move beyond isolated point solutions toward broad, interoperable models capable of understanding fusion plasmas in a comprehensive way.
@@
Progress toward either specialized or generalist plasma models is hampered by the absence of open, standardized benchmarks. Fusion datasets remain fragmented across institutions, locked behind proprietary interfaces, or stored in domain‑specific formats that are difficult for Machine Learning researchers to access or interpret (Strand et al., 2022). Without unified task definitions, metrics, or evaluation protocols, it becomes impossible to provide a fair comparison of methods and to measure progress systematically. A benchmark is therefore needed to frame the plasma modeling problems in a broader and more structured way, enabling cross‑comparison across algorithms, reproducibility across labs, and accessibility for researchers both inside and outside the fusion community.
At the same time, fusion diagnostic data provide a unique set of challenges going beyond those found in standard datasets from more established areas of AI. The data is inherently multi‑modal, multi‑rate, and multi‑dimensional, with wide variations in signal quality, noise characteristics, and information content (see Figure 1 for an illustrative example). Many diagnostics operate asynchronously, contain gaps, or suffer partial failures; others produce high‑dimensional structured outputs such as flux maps or spectrograms. This complexity demands dedicated modeling approaches that can handle missing data, fuse heterogeneous signals, different modalities, and reason across fast and slow timescales—capabilities that are difficult to stress‑test without a standardized benchmark.
@@ 1.2. TokaMark Benchmark Overview
To fill the need for comprehensive fusion plasma benchmarks, we introduce TokaMark, the first large, open benchmark for evaluating AI models trained on real fusion data.
Data. To the best of our knowledge, FAIR-MAST data represents the only openly available dataset of real tokamak diagnostics. Recent releases (Jackson et al., 2025, 2024) curate a collection of real experiments and corresponding diagnostics measurements from MAST tokamak. From FAIR-MAST, we select 39 signals across heterogeneous modalities, harmonize metadata and build standardized loaders.
Tasks. For TokaMark, we defined a diverse suite of 14 downstream tasks organized into 4 groups, designed to probe core capabilities required by AI models for fusion plasmas: (i) representation learning from heterogeneous diagnostics; (ii) temporal reasoning across fast and slow timescales; (iii) robustness to incomplete state information; and (iv) generalization across operating regimes. Rather than optimizing for a single downstream objective, the tasks span a cascade of physical processes—from fast magnetic response to slower, transport-driven evolution and long-horizon precursors of MHD activity—while remaining closely aligned with routine experimental workflows. Wherever possible, the tasks minimize reliance on expert-labeled targets, supporting self-supervised and weakly supervised formulations, and enabling systematic evaluation of transferable plasma representations.
Evaluation. We introduce a hierarchical evaluation protocol aligned with the structure of FAIR-MAST and with scientific objectives. This hierarchy assesses both low-level prediction quality and high-level scientific utility. For specialized models (single-task training), the hierarchy yields granular, signal-level diagnostics; for generalist models (FM-style pretrained and then fine-tuned across tasks), it provides modular comparisons across tasks and broad scientific objectives.
Baseline. Finally, we provide a strong yet accessible baseline: a multi-branch convolutional encoder–decoder architecture inspired by the previous works (Seo et al., 2023, 2024) and largely expanded to ingest heterogeneous inputs and predict our task-specific targets. The architecture is trained independently for each task.
@@
  • Benchmark design. We define 14 tasks organized into four groups, with a hierarchical evaluation protocol, standardized windowing, and error metrics.
  • Data packaging. We freeze a stable subset of FAIR-MAST data, resolve schema inconsistencies, standardize metadata and units, and release it for reproducibility and long-time compatibility.
  • Tools and API. We provide a Python package for task-specific data loading, processing and batching, masking and alignment utilities, and evaluation logic, integrated with the PyTorch stack.
  • Baseline models. We release a multi-branch convolutional encoder–decoder baseline with reference configurations and training scripts, establishing reproducible baselines across all tasks.
@@
Magnetics Flux loops, pickup coils; saddle coils Diagnostic 5 kHz; 50 kHz Profile
Kinetics Thomson scattering; interferometer Diagnostic 0.2 kHz; 4 kHz Profile; time series
@@
Fast magnetics Mirnov coils Diagnostic 500 kHz Profile
Currents Poloidal field coil currents; solenoid current, plasma current Diagnostic 4 kHz Profile; time series
@@
Equilibrium Shape parameters, $J_{\text{tor}}$ metrics; plasma boundary; flux map Derived 0.2 kHz Time series; profile; video
2. Preliminaries
To ground TokaMark benchmark in the reality of complex heterogeneous fusion data, we first introduce a signal taxonomy that organizes these measurements by function and modality, which allows us to provide a consistent vocabulary for describing inputs, targets, and tasks in the benchmark.
2.1. Data Taxonomy
@@
MAST is a spherical tokamak located in Culham, Oxfordshire (UK) and was operated by UKAEA and EURATOM from 1999 to 2013 (Sykes et al., 2001; Counsell et al., 2005; Meyer et al., 2009). MAST, and tokamaks in general, operates in short experimental cycles known as discharges or shots. The length of those cycles depends on the device size, and in the case of MAST typically last around 2–3 seconds. Over its operational lifetime, MAST produced more than 30,000 shots, with each shot containing diagnostic signals measuring various properties of the plasma, including the magnetic field, plasma temperature, shape parameters, and applied heating, to name a few. Recent works (Jackson et al., 2025, 2024) have created an open dataset of diagnostic data from a subset of the history of MAST. FAIR-MAST data contains 11,573 shots from the last five experimental campaigns on MAST. In this work, we utilized a total of 39 signals from FAIR-MAST data for the design of our benchmark tasks.
@@
We organize all the selected signals along several complementary axes, as summarized in Table 1. First, we group signals by category, which encodes their physical semantics—that is, the physical quantities they represent and how they are used by MAST workflows. This information corresponds to the first two columns of Table 1. Note that the second column contains both individual signals and subcategories. These intermediate groupings are introduced for brevity: they consolidate multiple signals that are typically used together (at least within the proposed benchmark) and share the same structural and functional properties. For example, the shape parameters subcategory includes several attributes jointly describing plasma geometry. A full list of the signals used in this work is provided in the appendix in Table A.1.
@@
Second, we distinguish signals by their origin, for which we define three classes: (i) diagnostics, corresponding to direct hardware measurements; (ii) actuators, representing controllable machine parameters used to steer plasma behavior; and (iii) derived signals, comprising quantities produced by reconstruction pipelines such as EFIT (Appel and Lupelli, 2018) (e.g., shape parameters or flux maps).
@@
The third axis in Table 1 is frequency, reflecting the wide range of sampling rates across diagnostics, spanning from 0.2 kHz up to 500 kHz. Finally, the fourth axis is modality, describing the structural form of each signal: (i) time series, represented as 1D tensors (scalar value over time); (ii) profiles, represented as 2D tensors (vector over time); and (iii) videos, represented as 3D tensors (image over time). These latter two axes are particularly important for guiding the design of AI model architectures and loss functions.
2.2. Structural Taxonomy of Tasks